# Spot the Fake: Large Multimodal Model-Based Synthetic Image Detection with Artifact Explanation

### v1

@@ Abstract
With the rapid advancement of Artificial Intelligence Generated Content (AIGC) technologies, synthetic images have become increasingly prevalent in everyday life, posing new challenges for authenticity assessment and detection. Despite the effectiveness of existing methods in evaluating image authenticity and locating forgeries, these approaches often lack human interpretability and do not fully address the growing complexity of synthetic data. To tackle these challenges, we introduce FakeVLM, a specialized large multimodal model designed for both general synthetic image and DeepFake detection tasks. FakeVLM not only excels in distinguishing real from fake images but also provides clear, natural language explanations for image artifacts, enhancing interpretability. Additionally, we present FakeClue, a comprehensive dataset containing over 100,000 images across seven categories, annotated with fine-grained artifact clues in natural language. FakeVLM demonstrates performance comparable to expert models while eliminating the need for additional classifiers, making it a robust solution for synthetic data detection. Extensive evaluations across multiple datasets confirm the superiority of FakeVLM in both authenticity classification and artifact explanation tasks, setting a new benchmark for synthetic image detection. The dataset and code will be released in: https://github.com/opendatalab/FakeVLM.
@@ 1 Introduction
As AI-generated content technologies advance, synthetic images are increasingly integrated into our daily lives [36, 11, 1]. Technologies like Diffusion [15, 36, 59] and Flux can generate highly realistic images. However, synthetic image data also poses significant risks, including potential misuse and social disruption [9, 20]. For instance, synthetic fake news [34, 19, 42] and forged facial scams [51, 31] make it increasingly challenging to discern the truth and establish trust in information. In response to these threats, the field of synthetic data detection has garnered widespread attention in recent years. However, traditional synthetic detection methods typically focus on simple authenticity judgments [2, 32, 56, 60, 39, 40], ultimately providing only a forgery probability or classification result. This approach still has significant limitations in terms of human interpretability. As a result, users find it challenging to understand the reasons behind the system’s decisions, affecting the decision-making process’s transparency and trustworthiness.
Figure 1: FakeVLM is a specialized large multimodal model designed for both DeepFake and general synthetic image detection tasks across multiple domains.
@@
The rapid development of large multimodal models has accelerated progress in synthetic data detection [52, 16, 24, 61, 17, 6]. These models, particularly, can provide explanations for authenticity judgments in natural language, thus laying the groundwork for enhancing the interpretability of synthetic detection. For instance, Jia et al. [18] explored ChatGPT’s ability to assess synthetic data, highlighting the potential of large models in this domain. LOKI [58] and Fakebench [28] further delved into the capabilities of large models in offering explanations for image detail artifacts. However, these studies primarily focus on pre-trained large models, utilizing strategies such as different prompt wordings to enhance model performance on this task rather than developing a specialized multimodal model for this specific domain. Moreover, existing general large models still show a significant performance gap compared to expert models or human users in detection tasks.
@@
Researchers have further explored and designed multimodal models specifically for synthetic data detection tasks [6, 17, 58]. For instance, works like DD-VQA [61] and FFAA [17] focus primarily on the performance of large models in Deepfake detection tasks, especially in the context of artifact explanation. However, their performance on more general types of synthetic images still requires further investigation. Other works, such as Fakeshield [52] and ForgeryGPT [24], effectively examine the ability of large models to localize forgery artifacts and explain manipulated synthetic data. However, artifacts in forged synthetic images often concentrate in transitional areas, exhibiting more noticeable edge artifacts. In contrast, direct synthetic images are more likely to show structural, distortion, or physical artifacts, highlighting significant differences between the two. Additionally, current large models still lag behind expert models in pure authenticity classification. For example, studies such as X2-DFD [6] and FFAA [17] attempt to integrate traditional expert-based synthetic detection methods to improve classification accuracy without fully unlocking the potential of large models in synthetic detection tasks.
@@
To address the challenges outlined above, we introduce FakeVLM, a large multimodal model specifically designed for fake image detection and artifact explanation. As shown in Fig 1, FakeVLM focuses on artifacts generated from synthetic image models rather than forgery artifacts. Moreover, it is not limited to facing Deepfake tasks but extends to more general synthetic data detection. Notably, FakeVLM achieves performance comparable to current expert models based on binary classification without requiring additional classifiers or expert models. Additionally, we introduce FakeClue, a dataset containing over 100,000 real and synthetic images, along with corresponding artifact cues for the synthetic images. FakeClue includes images from seven different categories (e.g., animal, human, object, scenery, satellite, document, deepfake), and leverages category-specific knowledge to annotate image artifacts in natural language using multiple LMMs. Our main contributions are as follows:
@@ 2.1 Synthetic Image Detection
Traditional synthetic detection tasks are typically approached as binary classification tasks [27, 18, 63, 14], treating them as either true or false, based on data-driven methods, with various detectors such as CNNs [22] and transformers [45]. Some research [5, 54, 44] has explored methods like learning anomalies in the frequency domain, reconstruction learning, and adversarial learning. The focus of traditional synthetic detection tasks has mainly been on addressing the generalization issue [32, 56], where the training and testing domains differ, in order to counter the rapidly evolving synthetic models. However, these methods only provide authenticity predictions without offering detailed explanations behind the predictions.
Some works have gone beyond the fundamental true/false classification problem, focusing on interpretable synthetic detection. For example, gradient-based methods are used to visualize highlighted regions of predictions [38, 41, 43]. Alternatively, model designs like DFGNN [21] enhance interpretability by applying interpretable GNNs to deepfake detection tasks. Additionally, research has explored the detection and localization of forgeries by constructing image artifacts or modifying labels [60, 39, 40]. While these methods enhance model interpretability, using natural language to describe the identified reasons remains underexplored.
@@ 2.2 Synthetic Image Detection via LMMs
Recently, the development of large multimodal models (LMMs) has been rapid. Models such as the closed-source GPT-4o [33] and open-source models like InternVL [7] and Qwen2-VL [46] have demonstrated outstanding performance across various tasks, showcasing impressive capabilities. In the domain of synthetic data detection, several works like [18], Fakebench [28], and LOKI [58] have investigated the potential of LMMs, where these large models not only deliver accurate synthetic detection judgments but also offer natural language explanations for their true/false predictions, enhancing the interpretability of synthetic detection. However, these studies mainly focus on evaluating pretrained large models rather than training expert multimodal models. Furthermore, existing general models still lag behind expert models or humans in detection tasks.
Figure 2: Construction pipeline of FakeClue dataset, including data collection from open source and self-synthesized datasets, pre-processing with categorization, label prompt design based on category knowledge, and multiple LMMs annotation with result aggregation.
An increasing number of studies have further explored the use of LMMs for synthetic detection, such as DD-VQA [61] and FFAA [17] have explored the performance of large models in the deepfake domain, yet the performance on more general synthetic image types remains underexplored. Fakeshield [52], SIDA [16] and ForgeryGPT [24] effectively investigate the ability of LMMs to detect and explain artifacts in manipulated synthetic data. However, tampering artifacts primarily focus on transitional artifacts, while direct image synthesis artifacts tend to involve structural distortions or other types of image warping, which differ significantly. Additionally, existing large models still lag behind expert models in simple real/false classification tasks [6, 18]. For example, studies like X2-DFD [6] and FFAA [17] have combined traditional expert synthetic detection methods with large models to improve classification accuracy.
@@ 3.1 Overview
We introduce FakeClue, a multimodal synthetic data detection benchmark designed for both general and DeepFake detection tasks. It includes two tasks: synthetic detection and artifact explanation, where the model is required to not only determine the authenticity of an image but also explain its artifacts. FakeClue covers 7 different categories of images, with over 100k image samples. Using a multi-LMM labeling strategy based on category priors, FakeClue is organized as image-caption pairs for both the image and its artifact explanation in natural language. Additionally, it focuses more on directly synthesized image artifacts rather than tampered image artifacts. The training and test sets are randomly split, with the test set containing 5,000 samples covering various image types. More detailed information about the dataset will be provided in the supplementary materials.
@@ 3.2 Construction of FakeClue
Data Collection: As shown in the data collection phase of Figure 2, FakeClue has two data sources, including open synthetic datasets and our newly synthesized data for specialized types. For open synthetic datasets, we extracted approximately 80K data from GenImage [64], FF++ [37] and Chameleon [55], maintaining a 1:1 ratio of fake to real data. For specialized types of data, such as remote sensing and document images, we generated the data ourselves. We collected remote sensing images from public datasets like CVUSA [49] and VIGOR [65], using GAN and Diffusion-based methods [57, 26], covering urban, suburban, and natural scenes. For document images, we adopted a layout-first, content-rendering approach, generating synthetic images of newspapers, papers, and magazines with real-world data sourced from the M6Doc dataset [8].
@@
Data Pre-Processing: At this stage, we first categorize the collected raw image data based on authenticity labels. Since the data from GenImage and Chameleon lack category information, we use a classification model [12] to divide the data into four categories: animals, objects, people, and landscapes. The FF++ dataset corresponds to the DeepFake category, while the newly synthesized satellite and document data also have clear category divisions. The distribution and proportions of these categories are shown in Figure 2. The labels obtained during the data preprocessing stage will serve as the foundation for the subsequent Label Prompt Design.
@@
Label Prompt Design based on category knowledge: To overcome the hallucinations and limited synthetic detection capabilities of large models, we inject external knowledge to aid in artifact detection. Based on the authenticity labels obtained in the data pre-processing phase, these labels are retained as prior knowledge within the prompt. For real images, we analyze the plausibility of the image as a photographic result, while for fake images, we focus on detecting artifacts throughout the image. Classification labels are used as category-specific knowledge to guide the model’s attention to key areas, such as the artifact detection cues for facial images, as shown in Fig. 2. Accordingly, we design 12 distinct types of prompts to address various authenticity and category labels. In addition, for synthetic images that may have high quality and no visible artifacts (e.g., images from the Chameleon dataset), we apply special image tags to allow the model to refrain from forced artifact explanation.
@@
Multiple LMMs Annotation: To mitigate the bias or hallucination effects of a single multimodal model, we adopt a strategy of annotating and then aggregating results from multiple high-performance open-source large models. For a given image set ${I_{i}}_{i=1}^{N}$, we first use three large multimodal models—Qwen2-VL, InterVL, and Deepseek—to generate candidate captions ${A_{i}^{1},A_{i}^{2},A_{i}^{3}}$ for each image. Different models may focus on different aspects of image artifacts in their captions. We then define an aggregation function $\mathcal{F}$ to combine these candidate captions into a unified annotation.
$A_{i}\;=\;\mathcal{M}\bigl{(}A_{i}^{1},A_{i}^{2},A_{i}^{3}\bigr{)}.$ (1)
@@
During this stage, the model extracts common points from multiple model responses, filters out irrelevant observations that appear in only one model (unless they are critical, like glaring artifacts) and organizes them into a hierarchical structure by categories such as texture, geometry, and lighting, before outputting in a fixed format.
Dataset Field Artifact Annotator Category Number
DD-VQA [61] DF Syn Human 5k
FF-VQA [17] DF Syn GPT 95K
LOKI [58] Gen Syn Human 3k
Fakebench [28] Gen Syn Human 6k
MMTD-Set [52] Gen Tam GPT 100k
FakeClue DF+Gen Syn Multi-LMMs 100k
@@ 3.3 Comparisons with Existing Datasets
Table 1 presents a comparison of the synthetic detection performance of FakeClue and existing evaluated LMMs datasets. FakeClue covers a broader range of domains, and unlike DD-VQA, FF-VQA is not limited to DeepFake detection tasks. The corresponding synthetic images exhibit well-defined category classifications, including specialized types such as satellite images and documents. In contrast, LOKI and Fakebench are limited by the number of annotations and can only serve as evaluation sets. Compared to the recent MMTD-Set dataset, FakeClue focuses more on directly synthesized image artifacts rather than tampered artifacts.

### v2

@@ Abstract
With the rapid advancement of Artificial Intelligence Generated Content (AIGC) technologies, synthetic images have become increasingly prevalent in everyday life, posing new challenges for authenticity assessment and detection. Despite the effectiveness of existing methods in evaluating image authenticity and locating forgeries, these approaches often lack human interpretability and do not fully address the growing complexity of synthetic data. To tackle these challenges, we introduce FakeVLM, a specialized large multimodal model designed for both general synthetic image and DeepFake detection tasks. FakeVLM not only excels in distinguishing real from fake images but also provides clear, natural language explanations for image artifacts, enhancing interpretability. Additionally, we present FakeClue, a comprehensive dataset containing over 100,000 images across seven categories, annotated with fine-grained artifact clues in natural language. FakeVLM demonstrates performance comparable to expert models while eliminating the need for additional classifiers, making it a robust solution for synthetic data detection. Extensive evaluations across multiple datasets confirm the superiority of FakeVLM in both authenticity classification and artifact explanation tasks, setting a new benchmark for synthetic image detection. The code, model weights, and dataset can be found here: https://github.com/opendatalab/FakeVLM.
@@ 1 Introduction
As AI-generated content technologies advance, synthetic images are increasingly integrated into our daily lives rombach2022high ; dhariwal2021diffusion ; abdullah2022chatgpt ; ye2025echo4oharnessingpowergpt4o . Technologies like Diffusion ho2020denoising ; rombach2022high ; zhang2023adding and Flux can generate highly realistic images. However, synthetic image data also poses significant risks, including potential misuse and social disruption cooke2024good ; ju2022fusing . For instance, synthetic fake news pawelec2022deepfakes ; jiang2023ai ; somepalli2023diffusion and forged facial scams xu2023combating ; mustak2023deepfakes make it increasingly challenging to discern the truth and establish trust in information. In response to these threats, the field of synthetic data detection has garnered widespread attention in recent years. However, traditional synthetic detection methods typically focus on simple authenticity judgments barni2020cnn ; ojha2023towards ; yandeepfakebench ; zhang2023perceptual ; shao2023detecting ; shao2024detecting , ultimately providing only a forgery probability or classification result. This approach still has significant limitations in terms of human interpretability. As a result, users find it challenging to understand the reasons behind the system’s decisions, affecting the decision-making process’s transparency and trustworthiness.
@@
The rapid development of large multimodal models has accelerated progress in synthetic data detection xu2024fakeshield ; huang2024sida ; li2024forgerygpt ; zhang2024common ; huang2024ffaa ; chen2024textit . These models, particularly, can provide explanations for authenticity judgments in natural language, thus laying the groundwork for enhancing the interpretability of synthetic detection. For instance, Jia et al. jia2024can explored ChatGPT’s ability to assess synthetic data, highlighting the potential of large models in this domain. LOKI ye2024loki and Fakebench li2024fakebench further delved into the capabilities of large models in offering explanations for image detail artifacts. However, these studies primarily focus on pre-trained large models, utilizing strategies such as different prompt wordings to enhance model performance on this task rather than developing a specialized multimodal model for this specific domain. Moreover, existing general large models still show a significant performance gap compared to expert models or human users in detection tasks.
@@
Researchers have further explored and designed multimodal models specifically for synthetic data detection tasks chen2024textit ; huang2024ffaa ; ye2024loki ; kang2025legionlearninggroundexplain ; yan2025gptimgevalcomprehensivebenchmarkdiagnosing . For instance, works like DD-VQA zhang2024common and FFAA huang2024ffaa focus primarily on the performance of large models in Deepfake detection tasks, especially in the context of artifact explanation. However, their performance on more general types of synthetic images still requires further investigation. Other works, such as Fakeshield xu2024fakeshield and ForgeryGPT li2024forgerygpt , effectively examine the ability of large models to localize forgery artifacts and explain manipulated synthetic data. However, artifacts in forged synthetic images often concentrate in transitional areas, exhibiting more noticeable edge artifacts. In contrast, direct synthetic images are more likely to show structural, distortion, or physical artifacts, highlighting significant differences between the two. Additionally, current large models still lag behind expert models in pure authenticity classification. For example, studies such as X2-DFD chen2024textit and FFAA huang2024ffaa attempt to integrate traditional expert-based synthetic detection methods to improve classification accuracy without fully unlocking the potential of large models in synthetic detection tasks.
@@
To address the challenges outlined above, we introduce FakeVLM, a large multimodal model specifically designed for fake image detection and artifact explanation. FakeVLM focuses on artifacts generated from synthetic image models rather than forgery artifacts. Moreover, it is not limited to facing Deepfake tasks but extends to more general synthetic data detection. Notably, FakeVLM achieves performance comparable to expert models based on binary classification without requiring additional classifiers or expert models. Additionally, we introduce FakeClue, a dataset containing over 100,000 real and synthetic images, along with corresponding artifact cues for the synthetic images. FakeClue includes images from seven different categories (e.g., animal, human, object, scenery, satellite, document, face manipulation), and leverages category-specific knowledge to annotate image artifacts in natural language using multiple LMMs. Our main contributions are as follows:
@@ 2.1 Synthetic Image Detection
Traditional synthetic detection tasks are typically approached as binary classification tasks li2018ictu ; jia2024can ; zhao2021learningselfconsistencydeepfakedetection ; haliassos2021lips , treating them as either true or false, based on data-driven methods, with various detectors such as CNNs lecun1998gradient and transformers vaswani2017attention . Some research chen2024learning ; yamada2025timeseriesanomalydetection ; tu2023hyperspectral has explored methods like learning anomalies in the frequency domain, reconstruction learning, and adversarial learning. The focus of traditional synthetic detection tasks has mainly been on addressing the generalization issue ojha2023towards ; yandeepfakebench , where the training and testing domains differ, in order to counter the rapidly evolving synthetic models. However, these methods only provide authenticity predictions without offering detailed explanations behind the predictions.
Some works have gone beyond the fundamental true/false classification problem, focusing on interpretable synthetic detection. For example, gradient-based methods are used to visualize highlighted regions of predictions selvaraju2017grad ; simonyan2013deep ; sundararajan2017axiomatic . Alternatively, model designs like DFGNN khalid2023dfgnn enhance interpretability by applying interpretable GNNs to deepfake detection tasks. Additionally, research has explored the detection and localization of forgeries by constructing image artifacts or modifying labels zhang2023perceptual ; shao2023detecting ; shao2024detecting . While these methods enhance model interpretability, using natural language to describe the identified reasons remains underexplored.
@@ 2.2 Synthetic Image Detection via LMMs
Recently, the development of large multimodal models (LMMs) has been rapid liu2023llava ; wen2025token ; liu2025shifting ; wen2025efficient ; wen2025stop . Models such as the closed-source GPT-4o gpt4o and open-source models like InternVL chen2024internvl and Qwen2-VL wang2024qwen2 have demonstrated outstanding performance across various tasks, showcasing impressive capabilities. In the domain of synthetic data detection, several works like jia2024can , Fakebench li2024fakebench , and LOKI ye2024loki have investigated the potential of LMMs, where these large models not only deliver accurate synthetic detection judgments but also offer natural language explanations for their true/false predictions, enhancing the interpretability of synthetic detection. However, these studies mainly focus on evaluating pretrained large models rather than training expert multimodal models. Furthermore, existing general models still lag behind expert models or humans in detection tasks.
Figure 1: Construction pipeline of FakeClue dataset, including data collection from open source and self-synthesized datasets, pre-processing with categorization, label prompt design based on category knowledge(Face M: Face manipulation), and multiple LMMs annotation with result aggregation.
An increasing number of studies have further explored the use of LMMs for synthetic detection, such as DD-VQA zhang2024common and FFAA huang2024ffaa have explored the performance of large models in the deepfake domain. However, their performance on general synthetic images remains less explored. Fakeshield xu2024fakeshield , SIDA huang2024sida and ForgeryGPT li2024forgerygpt investigate the ability of LMMs to detect/explain artifacts in manipulated synthetic data. However, tampering artifacts (mainly transitional), while direct image synthesis artifacts tend to involve structural distortions or other types of image warping, which differ significantly. Furthermore, current LMMs underperform expert models in simple real/false classification tasks chen2024textit ; jia2024can . For example, studies like X2-DFD chen2024textit and FFAA huang2024ffaa have combined traditional expert synthetic detection methods with large models to improve classification accuracy.
@@ 3.1 Overview
We introduce FakeClue, a multimodal synthetic data detection benchmark for general and DeepFake detection. It includes two tasks: synthetic detection and artifact explanation, requiring models to determine image authenticity and explain its artifacts. FakeClue covers 7 image categories with over 100k samples. Utilizing a multi-LMM labeling strategy with category priors, FakeClue provides image-caption pairs for the image and its natural language artifact explanation. It emphasizes direct synthesis over tampered image artifacts. Training/test sets are randomly split; the test set contains 5,000 diverse image samples. Detailed dataset information is provided in the supplementary materials.
@@ 3.2 Construction of FakeClue
Data Collection: As shown in the data collection phase of Figure 1, FakeClue has two data sources, including open synthetic datasets and our newly synthesized data for specialized types. For open synthetic datasets, we extracted approximately 80K data from GenImage zhu2024genimage , FF++ roessler2019faceforensicspp and Chameleon yan2024sanity , maintaining a 1:1 ratio of fake to real data. For specialized types of data, such as remote sensing and document images, we generated the data ourselves. We collected remote sensing images from public datasets workman2015localize ; zhu2021vigor ; ye2024cross ; zhou2025urbench ; ye2024crossview ; ye2025satellite , using GAN and Diffusion-based methods ye2024leveraging ; li2024crossviewdiff , covering urban, suburban, and natural scenes. For document images, we adopted a layout-first, content-rendering approach, generating synthetic images of newspapers, papers, and magazines with real-world data sourced from the M6Doc dataset Cheng_2023_CVPR .
@@
Data Pre-Processing: At this stage, we first categorize the collected raw image data based on authenticity labels. Since the data from GenImage and Chameleon lack category information, we use a classification model fang2024eva to divide the data into four categories: animal, object, human, and scene. The FF++ dataset corresponds to the face manipulation category, while the newly synthesized satellite and document data also have clear category divisions. The distribution and proportions of these categories are shown in Figure 1. The labels obtained during the data preprocessing stage will serve as the foundation for the subsequent Label Prompt Design.
@@
Label Prompt Design based on category knowledge: To overcome the hallucinations and limited synthetic detection capabilities of large models, we inject external knowledge to aid in artifact detection. Pre-processed authenticity labels serve as prior prompt knowledge. For real images, we analyze the plausibility of the image as a photographic result, while for fake images, we focus on detecting artifacts throughout the image. And classification labels are used as category-specific knowledge. This knowledge comprises predefined focal points and common artifact types pertinent to each category, guiding the model’s attention to key areas, such as the artifact detection cues for facial images, as shown in Fig. 1. Accordingly, we design 14 distinct types of prompts to address various authenticity and category labels. In addition, for synthetic images that may have high quality and no visible artifacts (e.g., images from the Chameleon dataset), we use special prompts to prevent the model from forcing interpretation (see Appendix H for detailed prompts for different categories).
@@
Multiple LMMs Annotation: To mitigate the bias or hallucination effects of a single multimodal model, we adopt a strategy of annotating and then aggregating results from multiple high-performance open-source large models. For a given image $I_{i}$ from the set $\{I_{i}\}_{i=1}^{N}$, we first use three large multimodal models—Qwen2-VL, InternVL, and Deepseek—to generate a set of candidate artifact captions $\mathcal{C}_{i}=\{A_{i}^{1},A_{i}^{2},A_{i}^{3}\}$. Each caption $A_{i}^{k}\in\mathcal{C}_{i}$ potentially highlights different aspects of image artifacts.
To synthesize a unified annotation $A_{i}$ from candidate captions $\mathcal{C}_{i}=\{A_{i}^{1},A_{i}^{2},A_{i}^{3}\}$, we employ Qwen2-VL for aggregation, denoted as $\mathcal{M}_{\text{agg}}$. This model is prompted with specific instructions $P_{\text{instr}}$ (detailed in the Appendix H) to identify consistent artifacts across $\mathcal{C}_{i}$ and remove redundant information. Thus, the final aggregated annotation $A_{i}$ for image $I_{i}$ is obtained as:
$A_{i}\;=\;\mathcal{M}_{\text{agg}}\bigl(\mathcal{C}_{i},P_{\text{instr}}\bigr)\quad$ (1)
Where $A_{i}^{k}$ is the $k$-th candidate caption for $I_{i}$, and $\mathcal{M}_{\text{agg}}$ aggregates $\mathcal{C}_{i}$ using prompt $P_{\text{instr}}$ (enforcing consistency/non-redundancy) into the final annotation $A_{i}$.
@@
The model extracts common points from multiple model responses, filters out irrelevant observations that appear in only one model (unless they are critical, like glaring artifacts) and organizes them hierarchically by categories (e.g., texture, geometry, lighting), before outputting in a fixed format.
@@
Dataset Field Artifact Annotator Category Number DD-VQA zhang2024common DF Syn Human 5k FF-VQA huang2024ffaa DF Syn GPT 95K LOKI ye2024loki Gen Syn Human 3k Fakebench li2024fakebench Gen Syn Human 6k MMTD-Set xu2024fakeshield Gen Tam GPT 100k FakeClue DF+Gen Syn Multi-LMMs 100k
@@ 3.3 Comparisons with Existing Datasets
Table 1 presents a comparison of the synthetic detection performance of FakeClue and existing evaluated LMMs datasets. FakeClue offers broader domain coverage and, unlike DD-VQA, is not confined to DeepFake detection, featuring well-defined categories including specialized types like satellite images and documents. In contrast, LOKI and Fakebench are limited by the number of annotations and can only serve as evaluation sets. Compared to the recent MMTD-Set dataset, FakeClue focuses more on directly synthesized image artifacts rather than tampered artifacts.